# The Pristine survey: XXVIII. The extremely metal-poor stream C-19 stretches over more than 100 degrees

### v1

@@ 3 Metallicity
For all the 15 stars observed in this work, we estimate the stellar parameters, the distance, and the extinction by fitting the metal-poor (, corresponding to ), -enhanced BASTI-IAC isochrones (Hidalgo et al., 2018; Pietrinferni et al., 2021) to the Gaia and 2MASS photometry (Cutri et al., 2003; Skrutskie et al., 2006) as well as the Gaia parallax of the star. We maximize the likelihood by varying the distance, initial mass, age, and extinction. We use a flat prior between 10 Gyr and 14 Gyr for the age and the initial mass function by Chabrier (2003) for the mass. We added error floors in quadrature to the reported uncertainties in photometry to account for possible systematic offsets between the observations and the isochrone models. The inferred stellar parameters are shown in Table 5.
@@
Combining all the follow-up programs, there are eleven stars that have high-resolution spectra with a resolving power 30 000 using Subaru/HDS, Magellan/MIKE, and VLT/UVES. Among them, eight stars have sufficiently high signal-to-noise ratio () spectra for the analysis of individual Fe lines (see in Tab. 4). The five stars with UVES spectra do not have a high enough for such analysis, but we are still able to infer metallicities from the Ca II triplet (CaT) lines for all of them. In order to check consistency between these two methods, we derive metallicities using the CaT lines for the nine stars with high- HR spectra. The two analysis methods are briefly described below.
@@
The first method is based on equivalent widths of individual Fe lines and stellar parameters. We first measure the equivalent widths by fitting a Gaussian profile to each Fe line. We then construct a model atmosphere by interpolating the grid of MARCS model atmospheres (Gustafsson et al., 2008), using the stellar parameters of the studied star. Since is not sufficient to determine the microturbulent velocity () for most of our targets, an empirical approach is adopted here. We first fit a linear function to the relation between and among the extremely metal-poor giants (, , and ) of Li et al. (2022a). The values we adopt for the C-19 stars are determined from this linear relation with the estimated . Based on these inputs (, , , and the equivalent widths), we derive the Fe abundance, using the radiative transfer code MOOG (Sneden, 1973), from 7–57 neutral Fe lines. The uncertainties on the derived metallicities are dominated by the uncertainties on and are estimated to be .
@@
In the second approach, we first measure the equivalent widths of the CaT lines by fitting a voigt profile to each component. We then derive the metallicity using the relation between metallicity, total equivalent width, and absolute band magnitude, as calibrated by Carrera et al. (2013). The metallicity uncertainties are dominated by the scatter around the calibrated relation (), yielding a typical uncertainty of in metallicity for this method.
@@
The metallicities reported in this work come from these two methods as well as from spectra with different resolutions. Thus, a consistency check is necessary. We first compare the metallicities of the nine stars with high HR spectra derived using the two methods described above. These stars are color coded by in the top panel of Fig. 3. By taking into account the uncertainties, the [Fe/H] metallicities based on the Fe lines and [Fe/H]CaT, derived from the CaT lines, are close to the one-to-one line. The most discrepant star (2640793013114783872) corresponds to the lowest spectra ( = 7 in Tab. 4), for which both methods may still work. Excluding this low- star, the overall comparison shows that metallicities derived from these two methods are consistent within dex. The metallicities derived from the Fe lines for the eight stars also agree well with the mean value of [Fe/H] = reported by Martin et al. (2022b), represented by the purple dashed lines and transparent bands in the top panel of Fig. 3.
@@
We then compare the metallicities for the same stars that are derived from spectra of different resolutions as well as from these two methods. There are three stars for which we can do this comparison. In the middle panel of Fig. 3, [Fe/H] represents the measurements from Fe lines using the different HR spectra while [Fe/H]CaT shows estimates using low-to-medium spectra from INT/IDS with 8 000, covering the CaT region. The different values for these three stars (yellow and magenta symbols) agree well with each other. Here, the two stars with VLT/UVES spectra are existing C-19 members, Pristine 355.32+27.59 (2865251577418971392) and Pristine 354.96+28.47 (2866151046649496832), from the C-19 discovery study (Martin et al., 2022b). The other star that has one Subaru/HDS spectrum and one INT/IDS spectrum is a new member confirmed in this study.
@@
To summarize this section, we selected sixteen candidate members of the C-19 stream based on their kinematics in Sec.2 and confirmed twelve new members that have spectroscopic metallicities consistent within the systematic uncertainty (0.2 dex) of the C-19 stream (Martin et al., 2022b) when we take into account the uncertainties of our methods. We plot the metallicity distributions (MDFs) of the newly confirmed members in the bottom panel of Fig. 3, where the blue and orange histograms denote the MDFs of [Fe/H] and [Fe/H]CaT respectively. There are three stars with metallicities beyond the systematic uncertainty of the C-19 stream, which are not included in the MDFs. These stars still all have low-metallicity values () and are marked as ‘probable members’ in Tab. 5, and highlighted by dotted circles in Fig. 2 and Fig. 4. We can see the probable members are not further away from the orbit than the confirmed ones, and they have similar confidence in membership shown in kinematic space. This may suggest that there was an additional, slightly more metal-rich population in the C-19 progenitor. Intriguingly, this result is similar to those of Bonifacio et al. (2024) for turnoff stars in the main body of C-19. However, these stars also all have low spectra and we cannot rule out that the low is unduly affecting our results. These stars require higher S/N spectra to confirm their metallicities.
4 Velocity Dispersion, Width, Mass
@@
Based on the full sample of 22 confirmed members with precise radial velocities, 10 stars from Yuan et al. (2022b) and 12 stars in Sec. 3, we are able to re-evaluate the dynamical properties of C-19. Here, we first rotate the equatorial coordinates to a set of stream coordinate (, ), following the technique presented in Koposov et al. (2010) and adopting the pole and zero-point of the main body of the stream (, , ) from Ibata et al. (2023). The resulting projection of the sample is displayed in Fig. 4, along with that of the stream orbit, from the same study, that is simultaneously fitted to 29 streams, for an isolated axisymmetric Milky Way potential (Ibata et al., 2023). We then calculate the offsets in velocity () and perpendicular angular distance () with respect to the orbit. As in Yuan et al. (2022b), we use the formalism of Martin et al. (2018) to infer the mean and dispersion of both quantities, and in this work we obtain shown in Fig. 5. Among all the members, the star above the disk has the largest offset from the orbit, as shown in the bottom panel of Fig. 4. By excluding this star, the velocity dispersion is only slightly smaller: . To estimate the stream width, we use the densest part of the stream, i.e., the main body located close to the apocentre with the distance around 20 kpc. The dispersion in angular distance is = , and the stream width is translated to be 200 pc. These new results clearly show that, over the 100∘ traced by the updated sample, the C-19 stream is even hotter than previously found. Previous studies focused on the main body of C-19 and gave a velocity dispersion of , a stream width of pc (Yuan et al., 2022b), and a velocity dispersion of for sub-giants (Bonifacio et al., 2024).
@@
The top panel of Fig. 4 shows that members brighter than (orange circles) are nearly evenly distributed along the orbit and become just slightly more clustered in the main body, where seven bright stars are located. There are now thirteen confirmed bright members (), which has doubled the previous number of known members presented by Yuan et al. (2022b). There is no reason to expect significant parts of the streams have been missed at this stage as stream segments with similar significance and similar orbital properties as C-19 should have been found using our search method in dynamical space by combining StarGO and STREAMFINDER. This is particularly true beyond the current extent of the stream, in the Galactic caps (Fig. 2), where it is easier to isolate streams because of their higher contrast with the Milky Way stellar populations. The one place where parts of the stream could still be evading detection is behind the Milky Way disk, north of the main body of C-19. The upcoming Gaia DR4 with deeper BP/RP spectra as well as further extensions of the Pristine footprint in this direction should help search for the presence of C-19 in this low-latitude region.
@@
With the current extension of the stream and the star-by-star search in Gaia, described in Sec.2.2, we have checked all the low-metallicity stars in the Pristine-Gaia synthetic catalogue along the stream orbit with compatible proper motions down to (Martin et al., 2024). Assuming the resulting sample of members is close to complete, we can update the estimated mass of the C-19 stream. This mass will still be a lower limit but should be closer to the true mass of the C-19 progenitor than earlier estimates ( or ) based on the main body of the stream (Martin et al., 2022a). By summing the fluxes of all the stars brighter than along the giant branch, the total luminosity is after correction for the unobserved fainter stars using the luminosity function associated with the PARSEC isochrone mentioned above (Bressan et al., 2012; Fu et al., 2018) and using a Kroupa initial mass function (IMF Kroupa, 2001, 2002). Assuming a mass-to-light ratio 2–3, typical of very old stellar populations (Maraston, 2005), the mass of C-19 is 3.7 – 5.5 , which, unsurprisingly given the much larger number of confirmed bright C-19 members, is significantly larger than the previous estimate. Note that a different IMF would also affect the derived mass-to-light ratio ( e.g., Cappellari et al., 2012). There is observational evidence suggesting that top-heavy IMFs are favored in extremely dense star-forming regions, such as the 30 Doradus star-forming region in the Large Magellanic Cloud (Schneider et al., 2018) or star burst galaxies at redshift 2–3 (Zhang et al., 2018; Guo et al., 2024). The C-19 progenitor was very likely a globular cluster, which could have been born in a similar environment of dense clouds and intense star-formation. Assuming a top-heavy IMF calibrated for very metal-poor GCs (Marks et al., 2012), the mass-to-light ratio further increases by a factor , which would result in an even larger progenitor mass of 5 – 8.
@@
As was discussed in Yuan et al. (2022b), the typical velocity dispersion of a GC stream is less than (Li et al., 2022b). However, with more spectroscopic data available for streams, more have been found to be dynamically hot. For example, the recent velocity dispersion of GD-1 is 7 – 29 (Malhan et al., 2019; Ibata et al., 2023), despite a progenitor mass of (Ibata et al., 2020). It has been suggested that this progenitor was accreted and pre-processed inside a dark matter subhalo (Carlberg, 2020; Malhan et al., 2021; Carlberg & Agler, 2023). In the case of C-19, the experiment of pre-processing a globular cluster in a dark matter subhalo remains a good explanation for the observed velocity dispersion (Errani et al., 2022). A recent study by Carlberg et al. (2024) shows the tidal heating of a cold globular cluster stream by CDM dark subhalos are able to increase the large velocity dispersion to values similar to the dispersion measured in the main body of C-19 (). But, the heated velocity dispersion values stay short of the velocity dispersion we find from the larger sample presented in this paper (). Other possible scenario to explain the high velocity dispersion would include heating by giant molecular clouds (see e.g., Amorisco et al., 2016) or, as recently showed from detailed N-body simulations of globular clusters (Wang, 2020; Gieles et al., 2021; Wang et al., 2024), the presence of black holes could enhance the relaxation and disruption of a GC, leading to a larger initial line-of-sight velocity dispersion. Born in a dense and extremely metal-poor region, the C-19 progenitor may also favor a top-heavy IMF, and the strong stellar feedback from over-abundant massive stars could cause the cluster to be born dynamically hot in the first place (Wang, 2020; Wang et al., 2021). Finally, unresolved binaries could also lead to an artificially large velocity dispersion. From the eight C-19 stars with radial velocity measurements from Gaia, there is no noticeable variation in velocities from our measurements and those from Gaia RVS (Katz et al., 2023). We do not see unexpectedly large uncertainties in each measurement either. Note that each star only has one precise velocity measurement from our follow-up and the mean value from Gaia DR3. Therefore, future multi-epoch spectroscopic observations will be crucial to quantify the contribution of binaries to the observed properties of C-19.
@@ 5 Conclusions
In this work, we searched for potential members of C-19 in dynamical space by applying StarGO to the low-significance stream stars from STREAMFINDER. With this technique, we found 4 new C-19 members, located and south from the main body. After Gaia DR3, we perform a star-by-star searches along the C-19 orbit using very low-metallicity star catalogues from the Pristine survey (both based on genuine Pristine data and on Gaia XP spectra; Martin et al. 2024). In total, we confirm twelve new members of the C-19 stream through high-resolution spectroscopy. All new members have velocities compatible with predictions from the orbit based on the main body of the stream (Martin et al., 2022a; Ibata et al., 2023), and their metallicities are, within the uncertainty of measurements, in agreement with the extremely metal-poor mean metallicity of the stream.
Our search for C-19 members in the bright regime () utilizes the all-sky Gaia data and we therefore expect that the distribution of these bright stars gives a relatively fair representation of the extent of the stream. Its main body, where it was originally detected, is centered around (Ibata et al., 2021; Martin et al., 2022b), and remains the densest part but we now show that it continuously extends down to in the South. A separate segment of the stream of four members is located at . One single member is found on the other side of the Milky Way disk, at , in the North. Overall, the C-19 stream is now shown to extend over on the sky. By summing up the fluxes from bright stars along the giant branch (), we revise the lower mass limit of the stream significantly upwards to 3.7 – 5.5 . Based on the velocity measurements of all members, accurate at the level, we revise upwards both the velocity dispersion of the stream to 11.1, and its Gaussian width to 200 pc. The stream is both hotter and more diffuse than previously reported (Martin et al., 2022a; Yuan et al., 2022b).
From a dynamical point of view, it remains a puzzle how to explain the large velocity dispersion with the relative small mass of the stream progenitor. Our suggested scenario of a preheating/disruption of C-19 cluster in its own dark-matter halo remains a possibility (Errani et al., 2022). In addition, tidal heating from the interaction of the stream with dark-matter subhalos (Carlberg et al., 2024) or giant molecular clouds are also possibilities. The scenario of a top-heavy IMF would result in an over-abundance of massive stars, whose stellar feedback could lead to a dynamically hotter initial globular cluster immediately after its birth. Finally, the contribution of binary systems to the observed velocity dispersion remains unknown and could significantly inflate it (Wang et al., 2024). It is crucial and timely to obtain multiple spectroscopic observations of known C-19 members to hunt for spectroscopic binaries if we are to understand the dynamical nature of this exceptional stream.

### v4

@@ 3 Metallicity
For all the 16 stars observed in this work, we estimated the stellar parameters, the distance, and the extinction by fitting the metal-poor (, corresponding to ), -enhanced BASTI-IAC isochrones (Hidalgo et al., 2018; Pietrinferni et al., 2021) to the Gaia and 2MASS photometry (Cutri et al., 2003; Skrutskie et al., 2006), as well as the Gaia parallax of the star. We maximized the likelihood by varying the distance, initial mass, age, and extinction. We used a flat prior between 10 Gyr and 14 Gyr for the age, and the initial mass function by Chabrier (2003) for the mass. We added error floors in quadrature to the reported uncertainties in photometry, to account for possible systematic offsets between the observations and the isochrone models. The inferred stellar parameters are shown in Table 5.
@@
Combining all the follow-up programs, there are 11 stars that have high-resolution spectra with a resolving power 30 000 using Subaru/HDS, Magellan/MIKE, and VLT/UVES. Among them, eight stars have sufficiently high signal-to-noise ratio (S/N) spectra for the analysis of individual Fe lines (see Tab. 4). The five stars with UVES spectra do not have a high enough S/N for such analysis, but we are still able to infer metallicities from the Ca II triplet (CaT) lines for all of them. In order to check consistency between these two methods, we derived metallicities using the CaT lines for the nine stars with high S/N HR spectra. The two analysis methods are briefly described below.
@@
The first method is based on equivalent widths of individual Fe lines and stellar parameters. We first measured the equivalent widths by fitting a Gaussian profile to each Fe line. We then constructed a model atmosphere by interpolating the grid of MARCS model atmospheres (Gustafsson et al., 2008), using the stellar parameters of the studied star. Since S/N is not sufficient to determine the microturbulent velocity () for most of our targets, an empirical approach was adopted here. We first fitted a linear function to the relation between and among the extremely metal-poor giants (, , and ) of Li et al. (2022a). The values we adopted for the C-19 stars were determined from this linear relation with the estimated . Based on these inputs (, , , and the equivalent widths), we derived the Fe abundance using the radiative transfer code MOOG (Sneden, 1973), from 7–57 neutral Fe lines. The uncertainties on the derived metallicities are dominated by the uncertainties on and are estimated to be .
@@
In the second approach, we first measured the equivalent widths of the CaT lines by fitting a Voigt profile to each component. We then derived the metallicity using the relation between metallicity, total equivalent width, and absolute band magnitude, as calibrated by Carrera et al. (2013). The metallicity uncertainties are dominated by the scatter around the calibrated relation (), yielding a typical uncertainty of in metallicity for this method.
@@
The metallicities reported in this work come from these two methods, as well as from spectra with different resolutions. Thus, a consistency check is necessary. We first compared the metallicities of the nine stars with high S/N HR spectra derived using the two methods described above. These stars are color coded by S/N in the top panel of Fig. 3. By taking into account the uncertainties, the [Fe/H] metallicities based on the Fe lines and [Fe/H]CaT, derived from the CaT lines, are close to the one-to-one line. The most discrepant star (2640793013114783872) corresponds to the lowest S/N spectra ( = 7 in Tab. 4), for which both methods may still work. Excluding this low S/N star, the overall comparison shows that metallicities derived from these two methods are consistent within dex.
@@
We then compared the metallicities for the same stars that are derived from spectra of different resolutions, as well as from these two methods. There are three stars for which we could do this comparison. In the middle panel of Fig. 3, [Fe/H] represents the measurements from Fe lines using the different HR spectra, while [Fe/H]CaT shows estimates using low-to-medium spectra from INT/IDS with 8 000, covering the CaT region. The different values for these three stars (yellow and magenta symbols) agree well with each other. Here, the two stars with VLT/UVES spectra are existing C-19 members Pristine 355.32+27.59 (2865251577418971392) and Pristine 354.96+28.47 (2866151046649496832), from the C-19 discovery study (Martin et al., 2022b). The other star that has one Subaru/HDS spectrum and one INT/IDS spectrum is a new member confirmed in this study.
@@
To summarize this section, we selected 17 candidate members of the C-19 stream based on their kinematics in Sec.2 and confirm 13 new members that have spectroscopic metallicities consistent within the systematic uncertainty (0.2 dex) of the C-19 stream (Martin et al., 2022b) when we take into account the uncertainties of our methods. We plot the metallicity distribution (MDF) of [FeI/H] from the C-19 members in Yuan et al. (2022b) as the purple histogram, as well as the MDFs of newly confirmed members, in the bottom panel of Fig. 3, where the blue and orange histograms denote the MDFs of [Fe/H] and [Fe/H]CaT, respectively. The mean of the old members [FeI/H] = 3.37, compared to the mean of the new members: [Fe/H] = 3.25 and [Fe/H]CaT = 3.22. The offset is less than the systematic uncertainty of 0.2 dex reported in Martin et al. (2022b). Assuming the metallicity distribution is Gaussian, and using a likelihood analysis similar to that of Martin et al. (2022b), the inference on the metallicity dispersion comes out as consistent with no dispersion for both sets of metallicities listed in Table 2. However, since the spectra of the majority of stars in this paper have rather low S/N, we postpone a detailed discussion on this topic to an upcoming paper (Venn et al., in prep) that reaches the same conclusion, but based on much higher S/N reobservations of several of the bright members presented here. There are three stars with metallicities beyond the systematic uncertainty of the C-19 stream which are not included in the MDFs. These stars still all have low metallicity values () and are marked as ”probable members” in Tab. 5 and highlighted by dotted circles in Fig. 2 and Fig. 4. One of them (2658115921889849472⋇) has a discrepancy in velocities derived from two bands due to the low-quality spectrum, as mentioned in Sec. 2. We can see that the probable members are not further away from the orbit than the confirmed ones and that they have a similar confidence in membership shown in kinematic space. This may suggest that there was an additional, slightly more metal-rich population in the C-19 progenitor. Intriguingly, this result is similar to those of Bonifacio et al. (2024) for turnoff stars in the main body of C-19. However, these stars also all have low S/N spectra and we cannot rule out that the low S/N is unduly affecting our results. Higher S/N spectra are required to confirm their metallicities.
4 Velocity dispersion, width and mass
@@
Based on the full sample of 23 confirmed members with precise radial velocities, 10 stars from Yuan et al. (2022b) and 13 stars in Sec. 3, we were able to reevaluate the dynamical properties of C-19. Here, we first rotated the equatorial coordinates to a set of stream coordinates (, ), following the technique presented in Koposov et al. (2010) and adopting the pole and zero point of the main body of the stream (, , ) from Ibata et al. (2023). The resulting projection of the sample is displayed in Fig. 4, along with that of the stream orbit from the same study, which is simultaneously fitted to 29 streams, for an isolated axisymmetric Milky Way potential (Ibata et al., 2023). We then calculated the offsets in velocity () and perpendicular angular distance () with respect to the orbit. As in Yuan et al. (2022b), we used the formalism of Martin et al. (2018) to infer the mean and dispersion of both quantities, and in this work we obtain , shown in Fig. 5. If we simply used a polynomial fitted orbit to the confirmed members, the dispersion is slightly smaller, . Among all the members, the star above the disk has the largest offset from the orbit, as shown in the bottom panel of Fig. 4. By excluding this star, the velocity dispersion is only slightly smaller: using the C-19 orbit, using a polynomial fitted orbit. To estimate the stream width, we used the densest part of the stream, i.e., the main body located close to the apocenter with the distance around 20 kpc. The dispersion in angular distance is = and the stream width is translated to be 200 pc. These new results clearly show that, over the 100∘ traced by the updated sample, the C-19 stream is even hotter than previously found. Previous studies focused on the main body of C-19 and gave a velocity dispersion of , a stream width of pc (Yuan et al., 2022b), and a velocity dispersion of for subgiants (Bonifacio et al., 2024).
@@
The top panel of Fig. 4 shows that members brighter than (orange circles) are nearly evenly distributed along the orbit and become just slightly more clustered in the main body, where seven bright stars are located. There are now 14 confirmed bright members (), which has doubled the previous number of known members presented by Yuan et al. (2022b). There is no reason to expect significant parts of the streams to have been missed at this stage, as stream segments with similar significance and similar orbital properties as C-19 should have been found using our search method in dynamical space by combining StarGO and STREAMFINDER. This is particularly true beyond the current extent of the stream, in the Galactic caps (Fig. 2), where it is easier to isolate streams because of their higher contrast with the Milky Way stellar populations. The one place where parts of the stream could still be evading detection is behind the Milky Way disk, north of the main body of C-19. The upcoming Gaia DR4 data release, with deeper BP/RP spectra, along with further extensions of the Pristine footprint in this direction, should help the search for the presence of C-19 in this low-latitude region.
@@
With the current extension of the stream and the star-by-star search in Gaia, described in Sec.2.2, we have checked all the low-metallicity stars in the Pristine-Gaia synthetic catalog along the stream orbit with compatible proper motions down to (Martin et al., 2024). Assuming the resulting sample of members is close to complete, we can update the estimated mass of the C-19 stream. This mass will still be a lower limit but should be closer to the true mass of the C-19 progenitor than earlier estimates ( or ) based on the main body of the stream (Martin et al., 2022a). By summing the fluxes of all the stars brighter than along the giant branch, the total luminosity is after correction for the unobserved fainter stars using the luminosity function associated with the PARSEC isochrone mentioned above (Bressan et al., 2012; Fu et al., 2018) and using a Kroupa initial mass function (IMF; Kroupa, 2001, 2002). Assuming a mass-to-light ratio 2–3, typical of very old stellar populations (Maraston, 2005), the mass of C-19 is 3.7 – 5.5 , which, unsurprisingly given the much larger number of confirmed bright C-19 members, is significantly larger than the previous estimate. We note that a different IMF would also affect the derived M/L ratio ( e.g., Cappellari et al., 2012). There is observational evidence suggesting that top-heavy IMFs are favored in extremely dense star-forming regions, such as the 30 Doradus star-forming region in the Large Magellanic Cloud (Schneider et al., 2018) or starburst galaxies at redshift 2–3 (Zhang et al., 2018; Guo et al., 2024). The C-19 progenitor was very likely a globular cluster, which could have been born in a similar environment of dense clouds and intense star formation. Assuming a top-heavy IMF calibrated for very metal-poor globular clusters (Marks et al., 2012), the M/L ratio further increases by a factor of , which would result in an even larger progenitor mass of 5 – 8.
@@
As discussed in Yuan et al. (2022b), the typical velocity dispersion of a GC stream is less than (Li et al., 2022b). However, with more spectroscopic data available for streams, more have been found to be dynamically hot. For example, the recent velocity dispersion of GD-1 is 7 – 29 (Malhan et al., 2019; Ibata et al., 2023), despite a progenitor mass of (Ibata et al., 2020). It has been suggested that this progenitor was accreted and preprocessed inside a dark matter subhalo (Carlberg, 2020; Malhan et al., 2021; Carlberg & Agler, 2023). In the case of C-19, the experiment of preprocessing a globular cluster in a dark matter subhalo remains a good explanation for the observed velocity dispersion (Errani et al., 2022). A recent study by Carlberg et al. (2024) shows the tidal heating of a cold globular cluster stream by CDM dark subhalos are able to increase the large velocity dispersion to values similar to the dispersion measured in the main body of C-19 (). But the heated velocity dispersion values stay short of the velocity dispersion we find from the larger sample presented in this paper (). Other possible scenarios to explain the high velocity dispersion would include heating by giant molecular clouds (see e.g., Amorisco et al., 2016) or, as recently shown from detailed N-body simulations of globular clusters (Wang, 2020; Gieles et al., 2021; Wang et al., 2024), the presence of black holes could enhance the relaxation and disruption of a GC, leading to a larger initial line-of-sight velocity dispersion. Born in a dense and extremely metal-poor region, the C-19 progenitor may also favor a top-heavy IMF, and the strong stellar feedback from overabundant massive stars could cause the cluster to be born dynamically hot in the first place (Wang, 2020; Wang et al., 2021). Finally, unresolved binaries could also lead to an artificially large velocity dispersion. From the eight C-19 stars with radial velocity measurements from Gaia, there is no noticeable variation in velocities from our measurements and those from Gaia RVS (Katz et al., 2023). We do not see unexpectedly large uncertainties in each measurement either. We note that each star only has one precise velocity measurement from our follow-up and the mean value from Gaia DR3. Therefore, future multi-epoch spectroscopic observations will be crucial for quantifying the contribution of binaries to the observed properties of C-19.
@@ 5 Conclusions
In this work, we searched for potential members of C-19 in dynamical space by applying StarGO to the low-significance stream stars from STREAMFINDER. With this technique, we found four new C-19 members, located and south from the main body. After Gaia DR3, we performed star-by-star searches along the C-19 orbit using very low-metallicity star catalogs from the Pristine survey (both based on genuine Pristine data and on Gaia XP spectra; Martin et al. 2024). In total, we confirm 12 new members of the C-19 stream through high-resolution spectroscopy. All new members have velocities compatible with predictions from the orbit based on the main body of the stream (Martin et al., 2022a; Ibata et al., 2023), and their metallicities are, within the uncertainty of measurements, in agreement with the extremely metal-poor mean metallicity of the stream.
Our search for C-19 members in the bright regime () utilizes the all-sky Gaia data and we therefore expect that the distribution of these bright stars gives a relatively fair representation of the extent of the stream. Its main body, where it was originally detected, is centered around (Ibata et al., 2021; Martin et al., 2022b), and remains the densest part; but we now show that it extends continuously down to in the south. A separate segment of the stream of four members is located at . One single member is found on the other side of the Milky Way disk, at , in the north. Overall, the C-19 stream is now shown to extend over on the sky. By summing up the fluxes from bright stars along the giant branch (), we revise the lower mass limit of the stream significantly upward to 3.7 – 5.5 . Based on the velocity measurements of all members, accurate at the level, we revise upward both the velocity dispersion of the stream to 10.9, and its Gaussian width to 200 pc. The stream is both hotter and more diffuse than previously reported (Martin et al., 2022a; Yuan et al., 2022b).
From a dynamical point of view, it remains a puzzle as to how to explain the large velocity dispersion with the relatively small mass of the stream progenitor. Our suggested scenario of a preheating and disruption of the C-19 cluster in its own dark matter halo remains a possibility (Errani et al., 2022). In addition, tidal heating from the interaction of the stream with dark matter subhalos (Carlberg et al., 2024) or giant molecular clouds are also possible. The scenario of a top-heavy IMF would result in an overabundance of massive stars whose stellar feedback could lead to a dynamically hotter initial globular cluster immediately after its birth. Finally, the contribution of binary systems to the observed velocity dispersion remains unknown and could significantly inflate it (Wang et al., 2024). It is crucial and timely to obtain multiple spectroscopic observations of known C-19 members to hunt for spectroscopic binaries if we are to understand the dynamical nature of this exceptional stream.
6 Data availability
The updated C-19 member list is provided at https://doi.org/10.5281/zenodo.15240247